# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ 3 Sample Recycling for Perturbation Selection
leads to an estimator of $B$. In this case, the estimator of $B$ has zero variance, and its bias tends to 0 as $h_{1},h_{2}\to 0$ because the vanishing rate of $o(h_{1}^{2})$ and $o(h_{2}^{2})$ surpasses that of $h_{1}^{2}$ and $h_{2}^{2}$, respectively. Likewise, $\sigma^{2}(\theta_{0})$ can be estimated by $2nh_{1}^{2}\mathrm{Var}\left[\widehat{\Delta}_{n,h_{1}}\right]$.
However, $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are typically unknown, and estimating them may be as challenging as estimating $\alpha^{\prime}(\theta_{0})$. Therefore, we turn to seek their surrogates. A straightforward approach is to use sample mean and variance, respectively. That is, we generate several sample versions of $\widehat{\Delta}_{n,h}$, denoted by $\left\{\widehat{\Delta}_{n,h}(l),l=1,2,...,p\right\}$, and then use their sample mean and variance as the surrogates for $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. Nevertheless, for any perturbation $h$, we need $np$ sample pairs [^4] Note that sometimes we also use “samples”. Throughout, the meanings of “samples” and “sample pairs” are same and we do not make distinctions. to obtain these surrogates, which is unavailable due to the limitation of sample size, especially when $n$ or $p$ is large. To address this limitation, we introduce a sample recycling technique, the bootstrap, to construct the sample mean and variance.
On the other hand, our primary objective is to estimate $\alpha^{\prime}(\theta_{0})$ rather than determining the optimal perturbation $h^{*}$. Consequently, the number of samples required for estimating $h^{*}$ should be relatively small compared to the total number of samples available. As a result, using bootstrap sample means as surrogates of the expectations in (10) may not provide sufficient accuracy to yield a relatively precise estimator of $B$ by solving the two equations in (10). To address this issue, we select $K$ different values of $h$, expanding (10) to encompass $K$ linear equations correspondingly. Then least-squares regression is applied to derive the slope of linear equations, i.e., the constant $B$.Similar procedure is also used to obtain the estimator of $\sigma^{2}(\theta_{0})$.
In the following, we will introduce the bootstrap technique for estimating the sample mean and variance and the least-squares regression procedure for estimating the constants, along with their theoretical guarantees.
@@ 3.1 Bootstrap sample mean and variance
In this subsection, we apply the bootstrap technique to establish surrogates of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ based on some pilot samples. We also provide their corresponding theoretical analyses.
The bootstrap technique is described as follows. For a fixed $h$, assume that $n_{b}$ sample pairs are generated, denoted by $\{(Y_{i}(\theta_{0}+h),Y_{i}(\theta_{0}-h)),i=1,2,...,n_{b}\}$. Correspondingly, as defined in (2), we obtain $n_{b}$ outputs $\bm{\Delta}(h)\triangleq\{\Delta_{i}(h),i=1,2,...,n_{b}\}$. We then select, with replacement, $n_{b}$ times from $\bm{\Delta}(h)$ independently and randomly to obtain a group of bootstrap samples $\bm{\Delta}^{*}(h)=\{\Delta_{i}^{*}(h),i=1,2,...,n_{b}\}$. Using these bootstrap samples, the estimator of $\alpha^{\prime}(\theta_{0})$ is expressed as
@@
It is worth mentioning that in this procedure, the bootstrap samples come from the same pilot samples (the $n_{b}$ sample pairs) and do not require the generation of extra samples. In essence, the bootstrap technique is based on the reuse of the pilot samples.
It is natural to expect that the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ will exhibit similar asymptotic properties as described in (5) and (6), respectively, when $n_{b}$ is sufficiently large. However, establishing these properties theoretically is not a straightforward task, because $\widehat{\Delta}^{b}_{n_{b},h}$ is derived from the resampling technique given the pilot samples. Therefore, these properties should be established based on the given pilot samples.
For the purposes described above, let $\mathbb{E}_{*}$ and $\mathrm{Var}_{*}$ denote the expectation and variance under the bootstrap probability measure $\mathbb{P}_{*}(\cdot)\triangleq\mathbb{P}(\cdot|\bm{\Delta}(h))$. Note that $\mathbb{P}_{*}$ is a probability conditional on pilot samples. Also, $\mathbb{P}_{*}$ is a discrete probability distribution, and its probability mass function is shown in Table 1.
Table 1: The value and probability of $\mathbb{P}_{*}$.
@@
The asymptotic properties of the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ are summarized in Theorem 1. Before proceeding, we make the following Assumption 3, which facilitates the investigation of the term $o(h^{2})$ in (5). Furthermore, this assumption aids in the analysis of the estimation error of $B$, which in turn assists in setting the initial perturbations.
@@
In many practical problems, the function $\alpha(\theta)$ is five times continuously differentiable. For example, in the $M/M/1$ queue, the average waiting time is a smooth function of the arrival and service rate (see Kleinrock and Gail, 1996). Now we provide the theorem, and the proof is provided in Appendix A.1.
@@
Theorem 1 indicates that by the resampling technique, we obtain the bootstrap samples whose mean and variance under $\mathbb{P}_{*}$ exhibit similar properties to those described in (5) and (6), respectively. However, it is worth mentioning the difference between Theorem 1 and (5) and (6). In particular, (11) eliminates the randomness from the bootstrap sampling in $\{\Delta_{1}(h),\Delta_{2}(h),...,\Delta_{n_{b}}(h)\}$, and (12) measures the degree of dispersion under $\mathbb{P}_{*}$. In contrast, (5) eliminates the randomness of simulating $\widehat{\Delta}_{n,h}$, and (6) measures the degree of dispersion of $\widehat{\Delta}_{n,h}$.
Although $n_{b}$ is the number of pilot samples and so it could be small compared with the number of total samples, the asymptotic results in Theorem 1 serve as a guide for illustrating the properties of $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$. Specifically, Theorem 1 demonstrates that $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ inherit the asymptotic properties of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. That is, they are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively. As a result, we can estimate $B$ and $\sigma^{2}(\theta_{0})$ through these linear relationships, which we will present in the next subsection.
@@ 3.2 Regression for perturbation selection
In our setting, we consider $K$ perturbation parameters $h_{1},h_{2},...,h_{K}$, where $K$ is finite. Using $Kn_{b}$ pilot samples, we construct a system of $K$ linear equations like (10) by (11) in Theorem 1, and we consider the vector version as follows.
@@ 3.2.1 Consistency of $\widehat{B}$
The following theorem presents the consistency of $\widehat{B}$, whose proof is provided in Appendix A.2. The conditions about $h_{k}(k=1,...,K)$ in Theorem 2 ensure that the regression error vanishes fast enough, which is necessary for the consistency of regression estimators. Note that we have assumed $h_{k}=c_{k}n_{b}^{\gamma}$ $(k=1,...,K)$ in the theorem, where $\gamma<0$, and $c_{k}\neq 0$ are constants. This choice of $h_{k}$ aims to minimize the second moment of the regression error $\mathcal{E}_{e}$ in (15).
@@
Suppose that Assumptions 1 and 3 hold. For any $k=1,...,K(K\geq 2)$, let $h_{k}=c_{k}n_{b}^{\gamma}$ $(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. Then,
@@
Theorem 2 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $-1/6<\gamma<0$, $\widehat{B}$ is a consistent estimator of $B$. Furthermore, from (18), the convergence rate of the bias of $\widehat{B}$ is $n_{b}^{2\gamma}$, and that of the variance of $\widehat{B}$ is $n_{b}^{-1-6\gamma}$. In addition, Theorem 2 provides valuable guidance for the selection of appropriate values of $h_{k}$ ($k=1,...,K$) during the pilot stage.
Except for fixed values of $c_{1},...,c_{K}$, they can be generated from a proper distribution $\mathcal{P}_{0}$. In this case, $H_{K}$ and $V_{K}$ are both random. Since $c_{1},...,c_{K}$ are coefficients in the perturbations, they must not equal 0, and are typically bounded. For these purposes, there are many optional distributions, such as the truncated normal distribution, denoted by $\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$. Specifically, for any $k=1,...,K$, let $c_{k}\stackrel{{\scriptstyle d}}{{=}}X\mathbbm{1}_{\{L\leq X\leq U\}}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ represents the same distribution, $X\sim\mathcal{N}\left(\mu_{0},\sigma_{0}^{2}\right)$, and $L>0$ and $U>0$ denote the lower bound and upper bound, respectively. Because $0<L\leq c_{k}\leq U$, $H_{K}$ and $V_{K}$ are bounded.
When $c_{1},...,c_{K}$ are generated from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ is
@@
where the last equality is because $c_{1},...,c_{K}$ are independent with the pilot samples and the bias result in Theorem 2. If $c_{k}\sim\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$ for $k=1,...,K$, then $c_{k}$’s are bounded. Therefore, in equation (19), $H_{K}$ and the constants in $o\left(n_{b}^{2\gamma}\right)$ are both bounded.
@@
Comparing (18) with (19) and (3.2.1), it becomes evident that when $c_{k}$’s are generated from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ remains $O\left(n_{b}^{2\gamma}\right)$, while the variance of $\widehat{B}$ increases by $O\left(n_{b}^{4\gamma}\right)$. However, it is important to note that the convergence rate of the MSE of $\widehat{B}$ remains unchanged. Therefore, during the pilot stage, the $h_{k}$’s selected based on the tradeoff between bias and variance in (18) maintain the same order as those selected using the tradeoff between bias and variance in (19) and (3.2.1).
Note that the first term of $\widehat{\bm{\beta}}_{e}$ is also an estimator of $\alpha^{\prime}(\theta_{0})$, and the second term is $\widehat{B}$. Therefore, similar analysis to the proof of Theorem 2 could yield the asymptotic properties of the first term. If the total simulation budget is used in the pilot stage, then the first term may seem like a reasonable estimator of $\alpha^{\prime}(\theta_{0})$. However, this estimator lacks efficiency, because all perturbations in the pilot stage are not taken at the optimal value.
Similar to Theorem 2, we provide Corollary 6 in Appendix A.3, which concludes the asymptotic property of $\widehat{\alpha}^{\prime}(\theta_{0})$.
@@ 3.2.2 Consistency of $\widehat{\beta}$v
To conduct a more precise analysis of the convergence of $\widehat{\bm{\beta}}_{v}$, additional information about $\sigma^{2}(\theta)$ is required. Hence, we provide the following Assumption 4.
@@
The rationality of Assumption 4 is provided in Remark 1. Under Assumption 4, we are able to refine (12) as follows,
@@
Now, we present Theorem 3 which demonstrates the convergence rate of the bias and variance of $\widehat{\bm{\beta}}_{v}$. The proof of Theorem 3 is provided in Appendix A.5.
@@
Suppose that Assumptions 2 and 4 hold. For any $k=1,...,K$ $(K\geq 1)$, let $h_{k}=c_{k}n_{b}^{\gamma}$ $(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. Then, as $n_{b}\to\infty$,
@@
Remark 1.
Assumption 4 serves the purpose of estimating the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ and substantiating the appropriateness of selecting $\gamma$. If Assumption 4 is violated, the non-random component of $\mathcal{E}_{v}$ transitions to $o(1)$ (refer to (17)), which could have a slower convergence rate than $O(h^{2})$ but still tends to 0. Meanwhile, the random component remains unaffected. For example, if $\theta_{0}=0$, $\sigma(\theta)=1+\sqrt{\theta}$ when $\theta>0$, $\sigma(\theta)=1-\sqrt{-\theta}$ when $-1/4<\theta\leq 0$ and $\sigma(\theta)=1/2$ when $\theta\leq-1/4$. Then, $\sigma^{\prime}(\theta_{0})$ is not bounded above and for any $40<h\leq 1/$ 4, it holds that $\sigma(\theta_{0}+h)=\sigma(\theta_{0})+\sqrt{h}$, $\sigma(\theta_{0}-h)=\sigma(\theta_{0})-\sqrt{h}$. Additionally, the expression for $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ is given by $(n_{b}-1)(\sigma^{2}(\theta_{0})+h)/(2n_{b}^{2}h^{2})+\phi(h)$. Therefore, for any $h_{1},...,h_{K}\in(0,1/4]$,
$\mathcal{E}_{v}=\left[\frac{(n_{b}-1)h_{1}}{2n_{b}^{2}h_{1}^{2}}+\phi(h_{1}),...,\frac{(n_{b}-1)h_{K}}{2n_{b}^{2}h_{K}^{2}}+\phi(h_{K})\right]^{\top}.$
Consequently, the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ is slower than what is illustrated in Theorem 3. At this point, the selection of $\gamma$ may cause confusion. This is because the convergence rate of $\mathrm{Bias}\left[\widehat{\bm{\beta}}_{v}\right]$ is slower than that of $\mathrm{Bias}\left[\widehat{B}\right]$, whereas the convergence rate of $\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ is faster than that of $\mathrm{Var}\left[\widehat{B}\right]$. In summary, Assumption 4 provides a greater rationale for the selection of $\gamma$.
Theorem 3 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $\gamma<0$, $\widehat{\bm{\beta}}_{v}$ is a consistent estimator of $\bm{\beta}_{v}$. The convergence rate of the variance of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{-1}$, which is independent of $\gamma$ and faster than that of $\widehat{B}$. Additionally, the convergence rate of the bias of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{2\gamma}$, which is equal to that of $\widehat{B}$. As a result, when comparing to $\widehat{B}$, $\widehat{\bm{\beta}}_{v}$ is easier to estimate.
3.3 The optimal perturbation in the pilot stage
From the asymptotic properties of $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ as shown in Theorems 2 and 3, respectively, we derive that the optimal perturbation in the pilot stage is not the same as the optimal perturbation $h^{*}$ of CFD problem as shown in (8). In fact, since $\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ is unrelated to $\gamma$, we only have to focus on (18) and choose a suitable $\gamma$ to balance the bias and variance of $\widehat{B}$. Specifically, according to Theorem 2, we consider the following optimization problem:

### v4

@@ 3 A Sample-driven Method for Constant Estimation
leads to an estimator of $B$. This estimator is noise-free and its deviation from the true value tends to 0 as $h_{1},h_{2}\to 0$ because the convergence rate of $o(h_{1}^{2})$ and $o(h_{2}^{2})$ is faster than that of $h_{1}^{2}$ and $h_{2}^{2}$, respectively. Likewise, $\sigma^{2}(\theta_{0})$ can be estimated by $2nh_{1}^{2}\mathrm{Var}\left[\widehat{\Delta}_{n,h_{1}}\right]$.
However, $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are typically unknown, and estimating them may be as challenging as estimating $\alpha^{\prime}(\theta_{0})$. Therefore, we turn to seek their surrogates. A straightforward approach is to use sample mean and variance, respectively. That is, we generate several sample versions of $\widehat{\Delta}_{n,h}$, denoted by $\left\{\widehat{\Delta}_{n,h}(l),l=1,2,...,p\right\}$, and then use their sample mean and variance as the surrogates of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. Nevertheless, for any perturbation $h$, we need $np$ sample pairsto obtain these surrogates, which is unavailable due to the limitation of sample size. To address this limitation, we introduce a resampling technique, the bootstrap, to construct the sample mean and variance.
On the other hand, our primary objective is to estimate $\alpha^{\prime}(\theta_{0})$ rather than determining the optimal perturbation $h^{*}$. Consequently, the number of samples used for estimating $h^{*}$ is relatively small compared to the total sample size. As a result, the bootstrap sample means as surrogates of the expectations in (10) may not provide sufficient accuracy to yield a relatively precise estimator of $B$ by solving the two equations in (10). To address this issue, we select $K$ different values of $h$, expanding (10) to encompass $K$ linear equations correspondingly. Then least-squares regression is applied to derive the slope of linear equations, i.e., the constant $B$. Similar procedure is also used to obtain the estimate of $\sigma^{2}(\theta_{0})$.
In the following, we introduce the sample-driven method, including the bootstrap technique for estimating sample mean and variance and the least-squares regression procedure for estimating the constants, along with its theoretical guarantees.
@@ 3.1 Bootstrap Sample Mean and Variance
In this subsection, we apply the bootstrap technique to establish surrogates of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ based on some pilot samples, and we provide their corresponding theoretical analyses.
The bootstrap technique is described as follows. For a fixed $h$, assume that $n_{b}$ sample pairs have been generated, denoted by $\{(Y_{i}(\theta_{0}+h),Y_{i}(\theta_{0}-h)),i=1,2,...,n_{b}\}$. Correspondingly, as defined in (2), we have $n_{b}$ outputs $\bm{\Delta}(h)\triangleq\{\Delta_{i}(h),i=1,2,...,n_{b}\}$. We pick with replacement $n_{b}$ times from $\bm{\Delta}(h)$ independently and randomly and then get a group of bootstrap samples $\bm{\Delta}^{*}(h)=\{\Delta_{i}^{*}(h),i=1,2,...,n_{b}\}$. Then, the estimator of $\alpha^{\prime}(\theta_{0})$ is expressed as
@@
It is worth mentioning that in this procedure, the bootstrap samples come from the same pilot samples (the $n_{b}$ sample pairs) and the generation of extra samples is not required. In essence, the bootstrap technique is based on the reuse of the pilot samples.
It is natural to expect that the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ will exhibit similar asymptotic properties as described in (5) and (6), respectively, when $n_{b}$ is sufficiently large. Let $\mathbb{E}_{*}$ and $\mathrm{Var}_{*}$ denote the expectation and variance under the bootstrap probability measure $\mathbb{P}_{*}(\cdot)\triangleq\mathbb{P}(\cdot|\bm{\Delta}(h))$. Note that $\mathbb{P}_{*}$ is a probability conditional on pilot samples. Also, $\mathbb{P}_{*}$ is a discrete probability distribution, and its probability mass function is shown in Table 1.
Table 1: The values and probabilities under $\mathbb{P}_{*}$.
@@
Asymptotic properties of the bootstrap mean and variance of $\widehat{\Delta}^{b}_{n_{b},h}$ are summarized in Theorem 1. Note that in the context of simulations, obtaining samples is typically expensive, while bootstrap operations are relatively cheap. Therefore, we adopt the analysis of bootstrap method from the field of statistics (Shao and Tu, 2012), i.e., theoretically studying $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ and using their Monte Carlo versions in practical applications.
Before going further, we make the following Assumption 3, which helps us examine the term $o(h^{2})$ in (5) and further assists us in analyzing the estimation error of $B$ and setting the initial perturbations for least-squares regression.
@@
In many practical problems, the function $\alpha(\theta)$ is five times continuously differentiable. For example, in the $M/M/1$ queue, the average waiting time is a smooth function of the arrival and service rate (see, Kleinrock and Gail, 1996). Now we provide the theorem and the proof is provided in Appendix A.1.
@@
Despite the relatively small size of $n_{b}$ compared to the total sample size, Theorem 1 provides valuable insights into the properties of $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$. Specifically, it demonstrates that $\mathbb{E}_{*}\widehat{\Delta}^{b}_{n_{b},h}$ and $\mathrm{Var}_{*}\left[\widehat{\Delta}^{b}_{n_{b},h}\right]$ inherit the asymptotic properties of $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$, respectively. That is, they are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively, and their slopes are respectively equal. As a result, we can estimate $B$ and $\sigma^{2}(\theta_{0})$ by leveraging these linear relationships, which we will detail in the following subsection.
@@ 3.2 Regression for Perturbation Selection
In our setting, we consider $K$ perturbation parameters $h_{1},h_{2},...,h_{K}$, where $K$ is finite. Using $Kn_{b}$ pilot samples and (11) in Theorem 1, we construct a system of $K$ linear equations like (10), and we consider the vector version as follows.
@@ 3.2.1 Consistency of $\widehat{B}$
The following theorem presents the consistency of $\widehat{B}$, whose proof is provided in Appendix A.2. The conditions about $h_{k}(k=1,...,K)$ in Theorem 2 ensure that the regression error vanishes fast enough, which is necessary for the consistency of regression estimators.
@@
Suppose that Assumptions 1 and 3 hold. For any $k=1,...,K(K\geq 2)$, denote $h_{k}=c_{k}n_{b}^{\gamma}(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, let $c_{j}\neq c_{k}$. Then,
@@
Theorem 2 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $-1/6<\gamma<0$, $\widehat{B}$ is a consistent estimator of $B$. Furthermore, from (18), the convergence rate of the bias of $\widehat{B}$ is $n_{b}^{2\gamma}$, and that of the variance of $\widehat{B}$ is $n_{b}^{-1-6\gamma}$. In addition, Theorem 2 provides a valuable guidance for the selection of appropriate values of $h_{k}$ ($k=1,...,K$) during the pilot stage.
Except for fixed values of $c_{1},...,c_{K}$, they can be generated from a proper distribution $\mathcal{P}_{0}$. In this case, $H_{K}$ and $V_{K}$ are both random. Since $c_{1},...,c_{K}$ are coefficients in the perturbations, they must not equal 0, and are typically bounded. For these purposes, there are many optional distributions, such as the truncated normal distribution, denoted by $\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$. Specifically, for any $k=1,...,K$, let $c_{k}\stackrel{{\scriptstyle d}}{{=}}X\mathbbm{1}\{L\leq X\leq U\}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ represents the same distribution, $X\sim\mathcal{N}\left(\mu_{0},\sigma_{0}^{2}\right)$, and $0<L<U$ denote the lower and upper bounds, respectively. Because $0<L\leq c_{k}\leq U$, $H_{K}$ and $V_{K}$ are bounded.
When we generate $c_{1},...,c_{K}$ from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ is
@@
where the last equality is because $c_{1},...,c_{K}$ are independent with the pilot samples and due to the bias result in Theorem 2. If $c_{k}\sim\psi\left(\mu_{0},\sigma_{0}^{2},L,U\right)$ for $k=1,...,K$, then $c_{k}$’s are bounded. Therefore, $H_{K}$ and the constants in $o\left(n_{b}^{2\gamma}\right)$ are all bounded.
@@
Comparing (18) with (19) and (3.2.1), it becomes evident that when $c_{k}$’s are generated from $\mathcal{P}_{0}$, the bias of $\widehat{B}$ still converges at a rate of $n_{b}^{2\gamma}$, while the variance of $\widehat{B}$ is increased by $n_{b}^{4\gamma}$. However, it is important to note that the convergence rate of the MSE of $\widehat{B}$ remains unchanged. Therefore, during the pilot stage, the $h_{k}$’s selected based on the tradeoff between bias and variance in (18) maintain the same order as those selected using the tradeoff between bias and variance in (19) and (3.2.1).
Note that when estimating $B$, the first element of the vector $\widehat{\bm{\beta}}_{e}$ is a rough estimator of $\alpha^{\prime}(\theta_{0})$, denoted by $\widehat{\alpha}^{\prime}(\theta_{0})$. In Appendix A.3, Corollary 6 establishes the consistency of $\widehat{\alpha}^{\prime}(\theta_{0})$. However, directly using $\widehat{\alpha}^{\prime}(\theta_{0})$ as the FD estimator is not recommended, as it is based on a small sample size of $Kn_{b}$. Despite this, it is evident that $\widehat{\alpha}^{\prime}(\theta_{0})$ contains valuable gradient information, which can be leveraged to construct the correlation-induced FD estimator in Section 4. Corollary 6 also plays a crucial role in proving the consistency of this new estimator.
@@ 3.2.2 Consistency of $\widehat{\beta}$v
To conduct a more precise analysis of the convergence of $\widehat{\bm{\beta}}_{v}$, we provide the following Assumption 4.
@@
Under Assumption 4, we are able to refine (12) as follows,
@@
Now we provide Theorem 3 which demonstrates the convergence rate of the bias and variance of $\widehat{\bm{\beta}}_{v}$. The proof of Theorem 3 is provided in Appendix A.5.
@@
Suppose that Assumptions 2 and 4 hold. For any $k=1,...,K$ $(K\geq 1)$, denote $h_{k}=c_{k}n_{b}^{\gamma}$ $(c_{k}\neq 0,\gamma<0)$ and for any $j\neq k$, $c_{j}\neq c_{k}$. Then, as $n_{b}\to\infty$,
@@
Theorem 3 indicates that for all $h_{k}=c_{k}n_{b}^{\gamma}$, where $k=1,...,K$ and $c_{k}\neq 0$, as long as $\gamma<0$, $\widehat{\bm{\beta}}_{v}$ is a consistent estimator of $\bm{\beta}_{v}$. The convergence rate of the variance of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{-1}$, which is independent of $\gamma$ and faster than that of $\widehat{B}$. Additionally, the convergence rate of the bias of $\widehat{\bm{\beta}}_{v}$ is $n_{b}^{2\gamma}$, which is equal to that of $\widehat{B}$.
3.3 The Optimal Perturbation for Pilot Samples
From the asymptotic properties of $\widehat{B}$ and $\widehat{\bm{\beta}}_{v}$ as shown in Theorems 2 and 3, respectively, we derive that the optimal perturbation the pilot samples is not the same as the optimal perturbation $h^{*}$ of CFD problem in (8). In fact, since $\mathrm{Var}\left[\widehat{\bm{\beta}}_{v}\right]$ is unrelated to $\gamma$, we only have to focus on (18) and choose a suitable $\gamma$ to balance the bias and variance of $\widehat{B}$. Specifically, according to Theorem 2, we consider the following optimization problem: